Please help

Imagine making a heavy door swing open by giving i
Answer #1

1)

Push 2 is more effective as the torque produced by the push is more in this case than the other

2)

Push 4 is more effective as the torque produced is 2 times the torque produced in Push 3

3)

Let the mass of each hanging mass be m

For the arrangement 1) (top-left one)

So, torque produced by the mass on the left, Tl = m*4d

Torque produced by the mass on the left,Tr = m*2d

As the two torques are not same . So they wont balance

For arrangement 2) (top-right)

Torque produced by the mass on left,Tl = m*4d = 4md

for the masses on the right,Tr = 2m*2d = 4md

As the torques are same, so they will balance out <------- balanced

For arrangement 3) (bottom left)

Tl = 3m*0.5d = 1.5*md

Tr = m*1.5d = 1.5md

So, it will balance out

Arrangement, 4)

Tl = 2m*4.5d = 9md

Tr = m*1.5d + m*2d +m*2.5d + m*3d = m*d*(1.5+2.5+2+3) = md*(4+5) = 9md

So, It will also balance .

B)

Here Tl = m*d + m*3.5*d = 4.5md

Tr = 3m*3d = 9md

So, the distance required = (9md - 4.5md)/(m) = 4.5*d